xx 11 22 33 44 55
P(x)P(x) 0.20.2 0.10.1 0.20.2 0.10.1 0.40.4

Given the discrete probability distribution above, determine the following:

(a) P(x>2)P(x>2) =

(b) P(x<5)P(x<5) =

(c) P(x≥4 or x<2)P(x≥4 or x<2) =
Answer #1

From given  probability distribution above,

(a)

P(x > 2) = 0.2 + 0.1 + 0.4 = 0.7

(b)

P(x < 5) = 0.2 + 0.1 + 0.2 + 0.1 = 0.6

(c)

(x≥4 or x<2) = 0.2 + 0.1 + 0.4 = 0.7